OpenStudy (anonymous):

solution for 4 m² = 4π R²

OpenStudy (anonymous):

\[R/m =\sqrt{\pi}\]

OpenStudy (anonymous):

Thanks, I should have clarified. Trying to solve for R but want to see the steps on how to do this.